OpenStudy (anonymous):

I am so confused. I have an exam tomorrow! Please help!!! I need a step by step explanation of these problems. Solve each equation for exact solutions in the interval [0, 2π]: a) 2sinx - √3 = 0 b) 2sinxcosx = √2 sinx c) 2sin2x = 1 - cosx

OpenStudy (anonymous):

\[2sinx-\sqrt{3} = 0 \rightarrow 2sinx=\sqrt{3} \rightarrow sinx=\frac{\sqrt{3}}{ 2 } \] \[sinx=\sin60^{0} \rightarrow x =60^{0}\]

OpenStudy (anonymous):

b. 2sinxcosx = √2 sinx \[ \frac{ sinx cosx }{ sinx } = \frac{ \sqrt{2} }{ 2 } \rightarrow cosx = \frac{ 1 }{ \sqrt{2} } \] \[cosx = \cos 45^{0} \rightarrow x = 45^{0}\]